# Self-similar inverse cascade from generalized symmetries

### v1

@@ Abstract
Despite the considerable attention these symmetries have received in vacuum (or equilibrium) physics, their role in non-equilibrium phenomena remains relatively unexplored. This gap in our understanding is particularly notable given the ubiquity of non-equilibrium processes in nature and their significance in various physical systems, from condensed matter physics, plasma physics, and astrophysics to cosmology.
In this Letter, we report a novel connection between higher-form symmetries and turbulent behavior, specifically focusing on the emergence of inverse cascade phenomena Kraichnan (1971); Fjørtoft (1953); Tabeling (2002); Alexakis and Biferale (2018). While inverse cascades have been well established in systems with conserved charges defined by integration over the entire space domain, such as energy and enstrophy in two-dimensional turbulence Kraichnan (1971); Fjørtoft (1953) or helicity in three-dimensional flows Moffatt (1969); Biferale et al. (2012), we demonstrate that higher-form symmetries, whose conserved charges are defined by integration over subspaces, can naturally induce a self-similar inverse cascade, where energy or conserved quantities flow hierarchically from small to large scales. We exemplify this mechanism for axion electrodynamics with the non-linear topological interaction, which possesses a 1-form symmetry and associated conservation law Sogabe and Yamamoto (2019); Hidaka et al. (2020, 2021b); Choi et al. (2023); Yokokura (2022). Our argument can also be applied to other systems with higher-form symmetries considered in, e.g., Ref. Yamamoto and Yokokura (2023). This suggests a new paradigm for understanding the emergence of coherent structures in turbulent systems from fundamental properties of higher-form symmetries, with potential implications across a broad range of physical scenarios.
@@
where $\phi$ is a dimensionless pseudo-scalar field identified as an axion, $f_{\mu\nu}=\partial_{\mu}a_{\nu}-\partial_{\nu}a_{\mu}$ is the field strength of a $U(1)$ gauge field $a_{\mu}$, $v$ is a parameter with mass dimension 1, $e$ is a coupling constant, and $C=1/(4\pi^{2})$. The symbol $\tilde{f}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}f_{\rho\sigma}$ with the totally anti-symmetric tensor $\epsilon^{\mu\nu\rho\sigma}$ denotes the dual of the field strength. The equations of motion of this system read
@@
$K_{a}({\cal S})=\frac{1}{2}\int_{\cal S}{\rm d}S^{\mu\nu}\frac{1}{e^{2}}\tilde{f}_{\mu\nu},\quad$ $T_{a}({\cal S})=-\frac{1}{2}\int_{\cal S}{\rm d}S^{\mu\nu}C\phi f_{\mu\nu}.$ (6)
with ${\cal V}$ and ${\cal S}$ being 3- and 2-dimensional closed subspaces. We have decomposed the conserved charges into those from the kinetic terms, $K_{\phi}({\cal V})$ and $K_{a}({\cal S})$, and those from the topological terms, $T_{\phi}({\cal V})$ and $T_{a}({\cal S})$. [^1] Note that $T_{\phi}({\cal V})$ and $T_{a}({\cal S})$ are not invariant under large gauge transformations. They can be made invariant under those transformations by rewriting them in terms of the topological quantum field theories (TQFT), and the resulting symmetries become non-invertible Choi et al. (2022); Cordova and Ohmori (2023); Choi et al. (2023); Yokokura (2022); Yamamoto and Yokokura (2023). However, as we will not consider the backgrounds with ’t Hooft lines or axionic vortices, this reformulation will not matter in the following discussion, and we will write these topological quantities without introducing TQFT.
@@
Dissipation.—In realistic situations in a medium, we must account for dissipative effects. As a concrete setup, we consider a system coupled to gapped charged matter, where the Ohmic current is absent. However, the axion field can still experience dissipation through its interactions with environmental degrees of freedom. By integrating out these modes, [^2] This procedure can be done systematically at the level of the action using the formulation of Refs. Crossley et al. (2017); Liu and Glorioso (2018). we obtain the following equation of motion for the axion:
@@
Set up for numerical simulations.—To demonstrate the emergence of self-similar inverse cascade, we numerically analyze the long-time behavior of the system. While the full $(3+1)$-dimensional dynamics involves complex interplay between all field components, the essential mechanism of generalized symmetry-induced cascade can be captured by studying field variations along a single direction. [^3] A similar approximation is employed in the analysis of the CPI in the presence of magnetic fields Hirono et al. (2015). Specifically, we assume the system is homogeneous along the $x^{2}$- and $x^{3}$-directions, and the field $a_{3}$ is spatially homogeneous, $a_{3}(x^{\mu})=a_{3}(x^{0})$. Under these conditions and in the temporal gauge $a_{0}=0$, the equations of motion reduce to
@@
In Fig. LABEL:fig:spec-functions, we plot the spectral functions $f_{\phi}(t,k)+f_{a_{2}}(t,k)$ and $g(t,k)$ of the energy and topological charge, respectively, at different times. The results show that the $a_{2}$ and $\phi$ fields, initially amplified by the generalized chiral instability, develop inverse cascades at later times.
@@
Both $N_{\rm kin}$ (or equivalently, $A(t)$) and $\xi^{-1}$ show asymptotic behavior $t^{-1/2}$, consistent with the predicted scaling (24), further validating our scaling analysis. It should be emphasized that this self-similar inverse cascade is due to the conservation law (7) associated with the 1-form symmetry and the transfer of the charge from the initial electric field to the topological charge.
@@
Summary and discussion.—In this Letter, we have demonstrated that higher-form symmetries can induce self-similar inverse cascade in a turbulent system. Using axion electrodynamics as a concrete example, we showed that starting from an initial electric field configuration that triggers the generalized chiral instability, the system exhibits characteristic scaling behavior at late times. Specifically, the spectral function $g(t,k)$ obeys the scaling law $g(lt,k)\sim lg(t,l^{1/2}k)$, while the instability parameter $A(t)$ decays as $t^{-1/2}$.
Interestingly, these scaling exponents match those of turbulence after the CPI Tashiro et al. (2012); Hirono et al. (2015); Yamamoto (2016); Buividovich and Ulybyshev (2016); Mace et al. (2020), suggesting a broader universality class encompassing systems with higher-form symmetries. This universality suggests similar phenomena may emerge in other systems with higher-form symmetries across diverse physical scenarios, from condensed matter to cosmology. The turbulent behavior we describe could be experimentally tested in magnetic materials where magnetic fluctuations serve as emergent axion fields Ooguri and Oshikawa (2012). Similar dynamics may also manifest in gauge field amplification during inflation Turner and Widrow (1988); Ratra (1992) in the presence of spectator axion-like fields.
@@
Acknowledgments.—This work was supported in part by JSPS KAKENHI Grant Numbers JP22H05111 (Y. H.), JP22H05111 (Y. H.), JP22H05118 (Y. H.), JP24K23186 (Y. H.), JP23K17687 (K. K.), JP22H01216 (N. Y.), JP24K00631 (N. Y.), JP21K13928 (R. Y.), JST, PRESTO Grant Number JPMJPR24K8 (Y. H.), and the National Natural Science Foundation of China (NSFC) under Grant Number 12347103 (K. K.).

### v2

@@ Abstract
Despite the considerable attention these symmetries have received in vacuum (or equilibrium) physics, their role in non-equilibrium phenomena remains relatively unexplored. Higher-form symmetries have provided valuable insights into the formulation of magnetohydrodynamics (MHD) Iqbal (2024) and analysis for the linear response of MHD Das et al. (2024), but their influence on strongly non-linear phenomena and turbulent behavior remains to be fully explored.
In this Letter, we report a novel connection between higher-form symmetries and turbulent behavior, [^1] We note that turbulence can arise in a wide range of systems governed by nonlinear field equations, not limited to hydrodynamic flows, with cascade phenomena serving as a hallmark of such turbulence. See, e.g., Ref. Berges et al. (2009) for studies of turbulence in non-Abelian gauge theories. The turbulence driven by the nonlinear topological interaction investigated in this work falls within this broader class. specifically focusing on the emergence of inverse cascade phenomena Kraichnan (1971); Fjørtoft (1953); Tabeling (2002); Alexakis and Biferale (2018). While inverse cascades have been well established in systems with conserved charges defined by integration over the entire space domain, such as energy and enstrophy in two-dimensional turbulence Kraichnan (1971); Fjørtoft (1953) or helicity in three-dimensional flows Moffatt (1969); Biferale et al. (2012), we demonstrate that higher-form symmetries, whose conserved charges are defined by integration over subspaces, can naturally induce a self-similar inverse cascade, where energy or conserved quantities flow hierarchically from small to large scales. We exemplify this mechanism for axion electrodynamics with the non-linear topological interaction, which possesses a 1-form symmetry and associated conservation law Sogabe and Yamamoto (2019); Hidaka et al. (2020, 2021b); Choi et al. (2023); Yokokura (2022). Our argument can also be applied to other systems with higher-form symmetries considered in, e.g., Ref. Yamamoto and Yokokura (2023). This suggests a new paradigm for understanding the emergence of coherent structures in turbulent systems from fundamental properties of higher-form symmetries, with potential implications across a broad range of physical scenarios.
@@
where $\phi$ is a $2\pi$ periodic pseudo-scalar field identified as an axion, $f_{\mu\nu}=\partial_{\mu}a_{\nu}-\partial_{\nu}a_{\mu}$ is the field strength of a $U(1)$ gauge field $a_{\mu}$, $v$ is a parameter with mass dimension 1, $e$ is a coupling constant, and $C=1/(4\pi^{2})$. The symbol $\tilde{f}^{\mu\nu}=\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}f_{\rho\sigma}$ with the totally anti-symmetric tensor $\epsilon^{\mu\nu\rho\sigma}$ denotes the dual of the field strength. In the following, we consider the dynamics of the system as a classical field theory. The equations of motion of this system read
@@
$K_{a}({\cal S})=\frac{1}{2}\int_{\cal S}{\rm d}S^{\mu\nu}\frac{1}{e^{2}}\tilde{f}_{\mu\nu},\quad$ $T_{a}({\cal S})=-\frac{1}{2}\int_{\cal S}{\rm d}S^{\mu\nu}C\phi f_{\mu\nu},$ (6)
with $\cal V$} and $\cal S$} being 3- and 2-dimensional closed subspaces. We have decomposed the conserved charges into those from the kinetic terms, $K_{\phi}({\cal V})$ and $K_{a}({\cal S})$, and those from the topological terms, $T_{\phi}({\cal V})$ and $T_{a}({\cal S})$. In the following, we assume that monopoles and axionic strings are sufficiently heavy so that they do not appear dynamically, ensuring that the associated conserved charges are well defined. [^2] It should be noted that, at the quantum level, $T_{\phi}({\cal V})$ and $T_{a}({\cal S})$ are not invariant under large gauge transformations. They can be made invariant under those transformations by rewriting them in terms of the topological quantum field theories (TQFT), and the resulting symmetries become non-invertible Choi et al. (2022); Cordova and Ohmori (2023); Choi et al. (2023); Yokokura (2022); Yamamoto and Yokokura (2023). However, as we will not consider the backgrounds with ’t Hooft lines or axionic vortices, this reformulation will not matter in the following discussion, and we will write these topological quantities without introducing TQFT. [^3] We also have higher-form symmetries associated with the Bianchi identities $\partial_{\mu}\tilde{f}^{\mu\nu}=0$ and $\epsilon^{\mu\nu\rho\sigma}\partial_{\rho}\partial_{\sigma}\phi=0$. Since these conservation laws do not involve the interaction terms, they do not contribute to the inverse cascade. Meanwhile, the presence of these symmetries implies the absence of dynamical magnetic monopoles and axionic strings.
@@
Dissipation.—In realistic situations in a medium, we must account for dissipative effects. As a concrete setup, we consider a system coupled to gapped charged matter, where the Ohmic current is absent. However, the axion field can still experience dissipation through its interactions with environmental degrees of freedom. Following the bottom-up approach to low-energy effective theory, we include the leading-order dissipative term allowed by symmetry in the axion sector. [^4] While a full treatment within the Schwinger-Keldysh effective field theory framework Crossley et al. (2017); Liu and Glorioso (2018) would clarify the associated Kubo-Martin-Schwinger constraints and confirm the absence of other symmetry-allowed terms at the same order, such an analysis is beyond the scope of the present work. We therefore treat this term as a phenomenological assumption consistent with symmetry. The corresponding low-energy equation of motion for the axion is given by
@@
(a) Spectral functions for energy and topological charge at different times. Both spectral functions exhibit the inverse cascade.
Set up for numerical simulations.—To demonstrate the emergence of self-similar inverse cascade, we numerically analyze the long-time behavior of the system. While the full $(3+1)$-dimensional dynamics involves complex interplay between all field components, the essential mechanism of generalized symmetry-induced cascade can be captured by studying field variations along a single direction. [^5] A similar approximation is employed in the analysis of the CPI in the presence of magnetic fields Hirono et al. (2015). Specifically, we assume the system is homogeneous along the $x^{2}$- and $x^{3}$-directions, and the field $a_{3}$ is spatially homogeneous, $a_{3}(x^{\mu})=a_{3}(x^{0})$. [^6] While this dimensional reduction simplifies the numerical analysis, we expect that the essential features of the inverse cascade should persist in the full (3+1)-dimensional case. This expectation is based on the general argument given earlier, where the interplay between dissipation and the conservation of the 1-form charge drives the transfer of the conserved quantity to larger spatial scales. This mechanism does not rely on the dimensional reduction and should apply in the full theory as well. Under these conditions and in the temporal gauge $a_{0}=0$, the equations of motion reduce to
@@
(b) Rescaled spectral functions for energy and topological charge at different times according to the scaling laws (26)–(28). Both rescaled spectral functions converge at later times, indicating self-similarity.
@@
In Fig. 2(a), we plot the spectral functions $f_{\phi}(t,k)+f_{a_{2}}(t,k)$ and $g(t,k)$ of the energy and topological charge, respectively, at different times. The results show that the $a_{2}$ and $\phi$ fields, initially amplified by the generalized chiral instability, develop inverse cascades at later times.
@@
Both $N_{\rm kin}$ (or equivalently, $A(t)$) and $\xi^{-1}$ show asymptotic behavior $t^{-1/2}$, consistent with the predicted scaling (24), further validating our scaling analysis.
@@
Summary and discussion.—In this Letter, we have demonstrated that higher-form symmetries can induce self-similar inverse cascade in a turbulent system. Using axion electrodynamics as a concrete example, we showed that starting from an initial electric field configuration that triggers the generalized chiral instability, the system exhibits characteristic scaling behavior at late times. In particular, the spectral function $g(t,k)$ grows according to the scaling law $g(lt,k)\sim lg(t,l^{1/2}k)$, while the instability parameter $A(t)$ decays as $t^{-1/2}$. The inverse cascade, in which spectral weight is transferred from short to long wavelengths, is ensured by the conservation law associated with the 1-form symmetry together with energy dissipation. To determine the scaling exponents, we imposed a self-similar ansatz and showed that the equations of motion admit such a solution. While the existence of a self-similar solution itself is non-trivial, it remains unclear a priori whether this solution is dynamically realized or acts as an attractor of the evolution. Our numerical simulations confirm that the system indeed approaches this self-similar scaling regime at late times.
Understanding the ultimate state of the system after the inverse-cascade scaling remains an open problem for future work. In the case of two-dimensional fluid turbulence in a finite system, which also exhibits an inverse cascade, it was initially conjectured that energy accumulates in low-energy modes, analogous to Bose-Einstein condensation Kraichnan (1967). This conjecture has since been examined in both numerical simulations and experiments Boffetta and Ecke (2012). A similar phenomenon may also be expected for the inverse-cascading turbulence studied here under finite-size conditions, although further investigation will be required.
The present results indicate the possible existence of a new universality class protected by higher-form symmetries. Interestingly, the scaling exponents we obtained coincide with those found in turbulence after the CPI Tashiro et al. (2012); Hirono et al. (2015); Yamamoto (2016); Buividovich and Ulybyshev (2016); Mace et al. (2020). This coincidence may suggest that universality classes protected by 0-form and 1-form symmetries are connected within a broader framework. Whether such a universality indeed exists, and how these connections are realized, remain open questions for future studies.
The phenomena studied here may also be realized in concrete physical systems. One promising class of candidates lies in condensed matter experiments, such as magnetic materials where magnetic fluctuations can act as emergent axion fields Ooguri and Oshikawa (2012). Another natural arena is cosmology. In particular, similar dynamics may manifest in gauge field amplification during inflation Turner and Widrow (1988); Ratra (1992) in the presence of spectator axion-like fields.
@@
Although we have not included monopoles and axionic strings in our analysis, it would be non-trivial how their existence affects the dynamics, since they would violate the conservation laws of $Q_{\phi}$, $Q_{a}$, and magnetic symmetries. Similar considerations also apply to the conventional CPI. We leave this problem for future study.
Acknowledgments.—This work was supported in part by JSPS KAKENHI Grant Numbers JP22H05111 (Y. H.), JP22H05118 (Y. H.), JP24K23186 (Y. H.), JP23K17687 (K. K.), JP22H01216 (N. Y.), JP24K00631 (N. Y.), JP21K13928 (R. Y.), JP25K17394 (R. Y.), JST, PRESTO Grant Number JPMJPR24K8 (Y. H.), and the National Natural Science Foundation of China (NSFC) under Grant Number 12547104 (K. K.).